# SocialFusion: Addressing Social Degradation in Pre-trained Vision-Language Models

### v1

@@ 4.1 Setup
Implementation details.  We train all SocialFusion models with the AdamW optimizer (loshchilov2018decoupled) using an initial learning rate of 2e42e-4, 500 warm-up steps, and cosine learning rate decay. We use a batch size of 32 for individual and joint training settings. Our LoRA adapters have a rank of 32; higher ranks did not improve performance in early experiments. For our visual-linguistic connector module, we use a hidden dimension of 4096. We experiment with three state-of-the-art VLMs: Qwen2-VL 2B (wang2024qwen2), Sail-VL 1.5 2B (dong-etal-2025-scalable), and MolmoE (deitke2024molmo), which is a mixture-of-experts model with 7 billion parameters in total and 1 billion active parameters. We use their recommended hyperparameters for training. Since these VLMs do not support native heatmap outputs, when making gaze predictions, we simply output the gaze points directly. This approach is viable for most VLMs because point data is included in their pre-training (deitke2024molmo; wang2024qwen2), unlike our SocialFusion, which would not be able to work with points in text without additional training in that domain.
Evaluation metrics.  For LAM, we report mAP and accuracy, as is done in the Ego4D paper (grauman2022ego4d). The HaGRIDv2 paper only reports mAP for gesture detection (nuzhdin2024hagridv2), which we report, but since many models we test achieve near-perfect mAP scores, we also report accuracy to provide another point of comparison. mollahosseini2017affectnet report several metrics for AffectNet, including AUCPR and accuracy. In this setting, AUCPR is simply a slightly different way to calculate mAP and yields almost identical results, so we also report mAP and accuracy for this dataset, remaining comparable to those baselines. In the case of PISC, following li2017dual, we report the mAP for each of the two classification subtasks: domain and relationship.
@@ 4.3 VLM Social Degradation Investigation Results
Compatibility.  We compute the cosine similarity between each task pair for each encoder at SocialFusion initialization and use these to calculate the GCD per encoder, which we report in Table 4. The results indicate that VLM pre-trained encoders slightly reduce inter-task alignment compared to the un-pre-trained versions. However, as noted by 10597963, gradients are considered conflicting when $GCD>1$. Since all tested encoders resulted in a $GCD<1$, the task gradients remain aligned in this setting. Therefore, the pre-training slightly reduces alignment, but the differences are small and the gradients are not pushed to direct conflict.
@@ 4.6 Comparison to the State of the Art
Finally, to better situate our method’s ability, we provide a comparison with the state of the art (SOTA) per task. We first focus on classification tasks. There have been many efforts to classify the social relationships in the two PISC subtasks (yu2024graph; li2022hf; li2020graph; zhang2019multi; goel2019end; wang2018deep). However, as seen in Table 7, SocialFusion significantly outperforms the SOTA approach (berlincioni2025navigating), highlighting the deep understanding that visual foundation models have of social scenes. On the test set, the SOTA is 87.0 mAP on PISC Domain and 79.5 on PISC Relations. The jointly-trained SocialFusion scores 94.2 mAP on PISC Domain and 88.5 mAP on PISC Relation, setting a new SOTA.
SocialFusion performs comparably or better than most baselines proposed in the original AffectNet paper (mollahosseini2017affectnet) but is not at the level of more recent SOTA methods (hassaballah2025integrating; ZHOU2025110110). HaGRIDv2 (nuzhdin2024hagridv2) is a recent extension over the HaGRID dataset (kapitanov2024hagrid), so we could not find prior work on it besides the baselines presented in the work’s paper. As shown in Table 7, the best baseline has an 89.4 mAP, which our jointly-trained model beats by a large margin with a mAP of 99.8 on the test set, setting the new state of the art. When it comes to LAM, there have also been several efforts dedicated to solving the task, with the recent state-of-the-art method using the InternVL image encoder and a Bi-LSTM network (lertniphonphan2024pcie_lam) to achieve a strong result. Our model performs at a competitive level, a few points behind this recent method.
Finally, on GazeFollow (recasens2015they), we achieve competitive results to prior work, shown in Table 1, but do not beat the state of the art (ryan2025gazelle) of 95.8 AUC, 0.041 minimum L2, and 0.099 average L2. Following ryan2025gazelle, we also attempted to train SocialFusion using DINOv2 and were able to nearly match the state of the art, achieving a 95.3 AUC, 0.048 minimum L2, and 0.105 average L2, making it likely that this DINO-based model can match the SOTA with some hyperparameter tuning. However, this version of SocialFusion performed worse on all other tasks, which shows that the choice of visual encoder can have significant impact on certain tasks. For reference, we include the full results using DINOv2 as the visual backbone in Appendix A.
@@ 5 Related Work
Visual social interaction understanding.  Building an intelligent system capable of understanding human social interaction is an underlying goal of many areas of research (li2025towards; recasens2015they; mollahosseini2017affectnet; li2018eye; lee2024modeling; kapitanov2024hagrid; ben2021video). Nonetheless, efforts toward classifying the various tasks required for visual social interaction understanding are very recent (lee2024socialaisurveyunderstanding; mathur2024advancing).
Notably, lee2024socialaisurveyunderstanding categorize visual social interaction tasks under two main categories: non-verbal social cue understanding and multimodal social cue understanding. These two categories are further split into six specific tasks. The former comprises gesture and body language analysis, gaze and attention analysis, and facial expression analysis, and the latter comprises multimodal emotion analysis, conversation dynamics analysis, and social situation analysis. This taxonomy lays the groundwork for building a robust, socially intelligent system, capable of analyzing visual and multimodal cues in a unified manner.
@@
Vision-language models for multitask learning.  Traditionally, methods in multitask learning (MTL) (zhang2021survey) train task-specific decoder heads or task-specific fine-tuned backbones to learn multiple tasks with some common parameters (kokkinos2017ubernet; xue2023egocentric). Since the development of strong general-purpose large language models and vision-language models (VLMs) (zhang2024vision), a new paradigm for MTL has emerged, wherein this large model solves multiple tasks in a unified manner (chen2023minigpt; shen2024multitask; ding2022prompt). VLMs can be repurposed on tasks unseen during training through in-context learning (ICL) (dong2022survey) or fine-tuning (zhang2024vision; xing2024survey). ICL can be done in a few-shot fashion (brown2020language), where a few examples of how to solve the task are included in the prompt to the model, or a zero-shot fashion, where the model is given no examples and instead relies only on the task description. Zero-shot learning is particularly difficult in the multi-modal case (cao2023review) due to VLMs’ task-agnostic training (zhang2024vision), especially with smaller VLMs on difficult tasks, and few-shot learning is not practically scalable with multiple datasets and long sequences of visual tokens in each example.
@@
Therefore, adapting the powerful pre-training of VLMs on downstream tasks through transfer learning has become common for difficult tasks (zhang2024vision; xing2024survey). Full or even partial direct fine-tuning, however, is often also impractical with these large models, giving rise to efficient methods like adapters (he2021effectiveness; xing2024survey), prefix-tuning (li2021prefix), prompt tuning (lester2021power; jia2022visual), low-rank adaptation (LoRA) (hu2022lora), and other variants (xing2024survey).
@@
The choice of a parameter-efficient fine-tuning method does not, by itself, resolve all underlying challenges of MTL. A key issue is negative transfer, where training on multiple tasks simultaneously degrades performance compared to training on tasks individually. This can stem from gradient conflict, where parameter updates for different tasks pull in opposing directions (zhang2021survey). Advanced MTL methods have been proposed to address this, such as PCGrad (yu2020gradient), which projects task gradients to remove any conflicting components, or other approaches that dynamically re-weight task losses or align gradients (chen2018gradnorm; liu2021towards). However, these methods primarily address optimization-level interference. While they could help slightly alleviate the problem due to the small increase in gradient conflict we observed in VLM visual encoders, they are not designed to solve the core issue identified in our work: a fundamental representation degradation. Thus, merely aligning gradients cannot fully address the social degradation in the pre-trained visual encoders.
@@
To provide a fair comparison, we compare our SocialFusion method with VLMs fine-tuned using a LoRA adapter, which is the same technique we use to fine-tune the underlying LLM behind SocialFusion. In particular, we experimented with three widely used VLMs of similar capacity: the 2B parameter variation of Qwen2-VL (wang2024qwen2), the 2B variation of Sail-VL 1.5 (dong-etal-2025-scalable), MolmoE (deitke2024molmo), a mixture-of-experts model with 1B active parameters and 7B parameters in total.
@@ Appendix B Negative transfer in InternVL3
Table 9: Like other VLMs, InternVL3 does not exhbit positive transfer in the joint setting.

### v2

@@ 4.1 Setup
Implementation details.  We train all SocialFusion models with the AdamW optimizer (Loshchilov and Hutter, 2019) using an initial learning rate of 2e42e-4, 500 warm-up steps, and cosine learning rate decay. We use a batch size of 32 for individual and joint training settings. Our LoRA adapters have a rank of 32; higher ranks did not improve performance in early experiments. For our visual-linguistic connector module, we use a hidden dimension of 4096. We experiment with three state-of-the-art VLMs: Qwen2-VL 2B (Wang et al., 2024), Sail-VL 1.5 2B (Dong et al., 2025), and MolmoE (Deitke et al., 2024), which is a mixture-of-experts model with 7 billion parameters in total and 1 billion active parameters. We use their recommended hyperparameters for training. Since these VLMs do not support native heatmap outputs, when making gaze predictions, we simply output the gaze points directly. This approach is viable for most VLMs because point data is included in their pre-training (Deitke et al., 2024; Wang et al., 2024), unlike our SocialFusion, which would not be able to work with points in text without additional training in that domain.
Evaluation metrics.  For LAM, we report mAP and accuracy, as is done in the Ego4D paper (Grauman et al., 2022). The HaGRIDv2 paper only reports mAP for gesture detection (Nuzhdin et al., 2024), which we report, but since many models we test achieve near-perfect mAP scores, we also report accuracy to provide another point of comparison. Mollahosseini et al. (2017) report several metrics for AffectNet, including AUCPR and accuracy. In this setting, AUCPR is simply a slightly different way to calculate mAP and yields almost identical results, so we also report mAP and accuracy for this dataset, remaining comparable to those baselines. In the case of PISC, following Li et al. (2017), we report the mAP for each of the two classification subtasks: domain and relationship.
@@ 4.3 VLM Social Degradation Investigation Results
Compatibility.  We compute the cosine similarity between each task pair for each encoder at SocialFusion initialization and use these to calculate the GCD per encoder, which we report in Table 4. The results indicate that VLM pre-trained encoders slightly reduce inter-task alignment compared to the un-pre-trained versions. However, as noted by Chai et al. (2024), gradients are considered conflicting when $GCD>1$. Since all tested encoders resulted in a $GCD<1$, the task gradients remain aligned in this setting. Therefore, the pre-training slightly reduces alignment, but the differences are small and the gradients are not pushed to direct conflict.
@@ 4.6 Comparison to the State of the Art
Finally, to better situate our method’s ability, we provide a comparison with the state of the art (SOTA) per task. We first focus on classification tasks. There have been many efforts to classify the social relationships in the two PISC subtasks (Yu et al., 2024; Li et al., 2022; 2020; Zhang et al., 2019; Goel et al., 2019; Wang et al., 2018). However, as seen in Table 7, SocialFusion significantly outperforms the SOTA approach (Berlincioni et al., 2025), highlighting the deep understanding that visual foundation models have of social scenes. On the test set, the SOTA is 87.0 mAP on PISC Domain and 79.5 on PISC Relations. The jointly-trained SocialFusion scores 94.2 mAP on PISC Domain and 88.5 mAP on PISC Relation, setting a new SOTA.
SocialFusion performs comparably or better than most baselines proposed in the original AffectNet paper (Mollahosseini et al., 2017) but is not at the level of more recent SOTA methods (Hassaballah et al., 2025; Zhou et al., 2025). HaGRIDv2 (Nuzhdin et al., 2024) is a recent extension over the HaGRID dataset (Kapitanov et al., 2024), so we could not find prior work on it besides the baselines presented in the work’s paper. As shown in Table 7, the best baseline has an 89.4 mAP, which our jointly-trained model beats by a large margin with a mAP of 99.8 on the test set, setting the new state of the art. When it comes to LAM, there have also been several efforts dedicated to solving the task, with the recent state-of-the-art method using the InternVL image encoder and a Bi-LSTM network (Lertniphonphan et al., 2024) to achieve a strong result. Our model performs at a competitive level, a few points behind this recent method.
Finally, on GazeFollow (Recasens et al., 2015), we achieve competitive results to prior work, shown in Table 1, but do not beat the state of the art (Ryan et al., 2025) of 95.8 AUC, 0.041 minimum L2, and 0.099 average L2. Following Ryan et al. (2025), we also attempted to train SocialFusion using DINOv2 and were able to nearly match the state of the art, achieving a 95.3 AUC, 0.048 minimum L2, and 0.105 average L2, making it likely that this DINO-based model can match the SOTA with some hyperparameter tuning. However, this version of SocialFusion performed worse on all other tasks, which shows that the choice of visual encoder can have significant impact on certain tasks. For reference, we include the full results using DINOv2 as the visual backbone in Appendix A.
@@ 5 Related Work
Visual social interaction understanding.  Building an intelligent system capable of understanding human social interaction is an underlying goal of many areas of research (Li et al., 2025; Recasens et al., 2015; Mollahosseini et al., 2017; Li et al., 2018; Lee et al., 2024a; Kapitanov et al., 2024; Ben et al., 2021). Nonetheless, efforts toward classifying the various tasks required for visual social interaction understanding are very recent (Lee et al., 2024b; Mathur et al., 2024).
Notably, Lee et al. (2024b) categorize visual social interaction tasks under two main categories: non-verbal social cue understanding and multimodal social cue understanding. These two categories are further split into six specific tasks. The former comprises gesture and body language analysis, gaze and attention analysis, and facial expression analysis, and the latter comprises multimodal emotion analysis, conversation dynamics analysis, and social situation analysis. This taxonomy lays the groundwork for building a robust, socially intelligent system, capable of analyzing visual and multimodal cues in a unified manner.
@@
Vision-language models for multitask learning.  Traditionally, methods in multitask learning (MTL) (Zhang and Yang, 2021) train task-specific decoder heads or task-specific fine-tuned backbones to learn multiple tasks with some common parameters (Kokkinos, 2017; Xue et al., 2023). Since the development of strong general-purpose large language models and vision-language models (VLMs) (Zhang et al., 2024), a new paradigm for MTL has emerged, wherein this large model solves multiple tasks in a unified manner (Chen et al., 2023; Shen et al., 2024; Ding et al., 2022). VLMs can be repurposed on tasks unseen during training through in-context learning (ICL) (Dong et al., 2022) or fine-tuning (Zhang et al., 2024; Xing et al., 2024). ICL can be done in a few-shot fashion (Brown et al., 2020), where a few examples of how to solve the task are included in the prompt to the model, or a zero-shot fashion, where the model is given no examples and instead relies only on the task description. Zero-shot learning is particularly difficult in the multi-modal case (Cao et al., 2023) due to VLMs’ task-agnostic training (Zhang et al., 2024), especially with smaller VLMs on difficult tasks, and few-shot learning is not practically scalable with multiple datasets and long sequences of visual tokens in each example.
@@
Therefore, adapting the powerful pre-training of VLMs on downstream tasks through transfer learning has become common for difficult tasks (Zhang et al., 2024; Xing et al., 2024). Full or even partial direct fine-tuning, however, is often also impractical with these large models, giving rise to efficient methods like adapters (He et al., 2021; Xing et al., 2024), prefix-tuning (Li and Liang, 2021), prompt tuning (Lester et al., 2021; Jia et al., 2022), low-rank adaptation (LoRA) (Hu et al., 2022), and other variants (Xing et al., 2024).
@@
The choice of a parameter-efficient fine-tuning method does not, by itself, resolve all underlying challenges of MTL. A key issue is negative transfer, where training on multiple tasks simultaneously degrades performance compared to training on tasks individually. This can stem from gradient conflict, where parameter updates for different tasks pull in opposing directions (Zhang and Yang, 2021). Advanced MTL methods have been proposed to address this, such as PCGrad (Yu et al., 2020), which projects task gradients to remove any conflicting components, or other approaches that dynamically re-weight task losses or align gradients (Chen et al., 2018; Liu et al., 2021). However, these methods primarily address optimization-level interference. While they could help slightly alleviate the problem due to the small increase in gradient conflict we observed in VLM visual encoders, they are not designed to solve the core issue identified in our work: a fundamental representation degradation. Thus, merely aligning gradients cannot fully address the social degradation in the pre-trained visual encoders.
@@
To provide a fair comparison, we compare our SocialFusion method with VLMs fine-tuned using a LoRA adapter, which is the same technique we use to fine-tune the underlying LLM behind SocialFusion. In particular, we experimented with three widely used VLMs of similar capacity: the 2B parameter variation of Qwen2-VL (Wang et al., 2024), the 2B variation of Sail-VL 1.5 (Dong et al., 2025), MolmoE (Deitke et al., 2024), a mixture-of-experts model with 1B active parameters and 7B parameters in total.
@@ Appendix B Negative transfer in InternVL3
Table 9: Like other VLMs, InternVL3 does not exhibit positive transfer in the joint setting.